OpenStudy (anonymous):

how can I delete indetermination in the following expression. Limx(x->0)((sqrt(x+2)-sqrt(2))/x)

OpenStudy (anonymous):

\[limit _{x->0}(\sqrt{x+2}-\sqrt{2})/2\] this is the equation

OpenStudy (anonymous):

conjugation

OpenStudy (anonymous):

limitx−>0(x+2−−−−−√−2−−√)/x Sorry

OpenStudy (anonymous):

you mean by multiplyting the numerator and the denominator by \[\sqrt{x+2}+\sqrt{2}\]

OpenStudy (anonymous):

Top and bottom

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

Changuanas, I actually get rid of indetermination. However when I apply the limit x->0, the answer is not the corrrect. The answer int he book is 1/2sqrt(2).

OpenStudy (anonymous):

Sorry. I was wrong. I already got the answer! Can you help me with another exercise? Just to check

OpenStudy (anonymous):

just post question, if fall asleep someone else would help

OpenStudy (anonymous):

\[Limit _{x->infinity} \sqrt{3x ^{4}-2x+4}+x\]

OpenStudy (anonymous):

I don't remember what happens to infinity under square root. Does it go to infinity?

OpenStudy (anonymous):

According to the graph, it increases with no bound

OpenStudy (anonymous):

Ignore all others and consider only highest exponent of x. What is highest exp?

OpenStudy (anonymous):

I'm sorry. Did you already have an answer: infinity? or it needs more work?

OpenStudy (anonymous):

yes, more work :)

OpenStudy (anonymous):

I think the answer is 1

OpenStudy (anonymous):

what is highest exponent x in your problem?

OpenStudy (anonymous):

inside the root? its 2

OpenStudy (anonymous):

Ignore all others divide sq rt 3x^4 by sq rt x^4

OpenStudy (anonymous):

Wouldn't that one have to be just infinity?

OpenStudy (anonymous):

I think he is in an early class. The class before you study infinity, right Emun?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In that case, don't ignore everything else, that is what you do in higher math;keep everything and divide each term by x^2

OpenStudy (anonymous):

It's funny being all done finals but still here during my break lol. why am I here?

OpenStudy (anonymous):

except that expression isn't anything you would need to do that for. maybe if you were dividing by a polynomial

OpenStudy (anonymous):

You here to help Emun I'm going to sleep.

OpenStudy (anonymous):

guys the equation was wrong. The equation is the following: \[\sqrt{x ^{2}-2x+4}+x\] I factored the expression inside the root out and I got \[\sqrt{(x-2)^{2}}+x\] so finally I got \[x-2+x\] as I apply the limit as x=infinity,the answer I got is 2

OpenStudy (anonymous):

Thanks changaunas :)

OpenStudy (anonymous):

The answer would be infinity, wouldn't it?

OpenStudy (anonymous):

wouldn't you get 2x-2 , which as x becomes large.. your expression does too? infinity!

OpenStudy (anonymous):

Yeah and this is the answer: Thanks gusy

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (\sqrt{x+2}-\sqrt{2})/x \times (\sqrt{x+2}+\sqrt{2})/(\sqrt{x+2}+\sqrt{2})\] \[\lim_{x \rightarrow 0} (x+2-2)/(x(\sqrt{x+2}+\sqrt{2}))\] \[\lim_{x \rightarrow 0} x/(x(\sqrt{x+2}+\sqrt{2}))\] \[\lim_{x \rightarrow 0} 1/(\sqrt{x+2}+\sqrt{2})\] \[1/(\sqrt{0+2}+\sqrt{2})\] \[1/(\sqrt{2}+\sqrt{2})\] \[1/(2\sqrt{2})\]